Session: Oct 2009 (session 3)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Yebes (Susana Garcia at Thu Oct 29 22:23:56 2009) Minor failures :
Because of technical problems, we started recording data at scan
number 169 at 07:04:42 DOY=301.
No other known problems.
Medicina (Roda at Thu Oct 29 22:28:48 2009) Minor failures : LATE
Technical problems, we started recording data at scan number 6
Onsala (Michael Lindqvist at Fri Oct 30 07:04:44 2009) Success :
* Some light rain.
* communication problem with the telescope (around 2009.302.10:07, see log) but I don't think we lost any data.
* A few bad scan_check, 2009.301.21:39:33.71/scan_check/138,ep066_on_no00442,?,,,,,
Effelsberg (Uwe Bach at Fri Oct 30 08:54:28 2009) Success :
Noto (P. Cassaro at Fri Oct 30 13:52:24 2009) Success :
No known problems
Jodrell_Bank (Alastair Gunn at Wed Nov 18 14:18:20 2009) Success :
The MkII telescope was used (Jb2).
No known problems.